# deepSURF: Detecting Memory Safety Vulnerabilities in Rust Through Fuzzing LLM-Augmented Harnesses

### v1

@@ Abstract
To address these limitations, we present deepSURF, a tool that integrates static analysis with Large Language Model (LLM)-guided fuzzing harness generation to effectively identify memory safety vulnerabilities in Rust libraries, specifically targeting unsafe code. deepSURF introduces a novel approach for handling generics by substituting them with custom types and generating tailored implementations for the required traits, enabling the fuzzer to simulate user-defined behaviors within the fuzzed library. Additionally, deepSURF employs LLMs to augment fuzzing harnesses dynamically, facilitating exploration of complex API interactions and significantly increasing the likelihood of exposing memory safety vulnerabilities. We evaluated deepSURF on 27 real-world Rust crates, successfully rediscovering 20 known memory safety bugs and uncovering 6 previously unknown vulnerabilities, demonstrating clear improvements over state-of-the-art tools.
@@ 1 Introduction
Rust’s focus on prioritizing memory safety while maintaining high performance makes it a compelling competitor in the systems programming space. Unlike traditional low-level languages like C/C++, which rely on manual memory management and are prone to common memory safety bugs such as buffer overflows, Rust’s unique set of safety rules eliminates these risks as early as during compilation. This feature has enabled its adoption in mainstream operating systems [21, 35], while its public registry of over 180k crates [^1] In Rust, a crate is a unit of code that can be a library or a binary. highlights its growing popularity  [12].
@@
Unsafe Rust undermines the language’s memory safety guarantees, introducing memory safety vulnerabilities. Microsoft and Google report that over 70% of critical bugs stem from memory safety issues [34, 47], leading agencies such as CISA to advocate for memory-safe programming [14, 27]. To date, around 27% of Rust bugs in the RustSec Advisory Database [46] are related to memory corruption, resulting from incorrect use of unsafe Rust [30, 49]. This significant percentage has motivated research into static[7, 31, 33] and dynamic [13, 5, 36] analysis tools to detect such issues.
@@
Static analyzers such as Rudra [7] and Yuga [31] aim to identify certain bug patterns in Rust source code. Although these tools successfully detect memory safety bugs, significant human involvement is required to filter their output (with high false positive rates) and to create proof-of-concept (PoC) test cases to confirm each bug. In contrast, dynamic analysis approaches, such as fuzzing, due to their context-sensitive analysis tend to have lower false positives and can uncover a broader range of bugs [36, 5, 39]. Although fuzzers are effective at detecting memory safety violations, their utility is often limited in the Rust ecosystem, where most code is provided as libraries rather than standalone binaries that can be directly tested by the fuzzer [23]. Consequently, before fuzzing Rust libraries, it is necessary to generate appropriate harnesses that encapsulate the library’s functionality and convert the fuzzer’s input into well-typed, valid data types expected by the target functions.
@@
Harness generation in Rust is tedious and time-consuming due to its complex type system and syntax. To ease this burden, several tools use static analysis to automate harness generation. However, despite improving test coverage, these tools struggle to detect memory safety bugs. RULF [23] and FRIES [51] lack support for traits, and RuMono [54] does not handle closures—both crucial for modeling buggy user-defined behavior. RPG [50] targets all unsafe code, including explicitly unsafe APIs, resulting in false positives. Additionally, these tools use heuristics and API dependency analysis to generate API call sequences. While the sequences can be syntactically valid, their API calls are often not semantically related and fail to reflect realistic usage patterns. Due to these limitations, although these tools have discovered bugs in various Rust crates, none of the reported issues have involved memory safety.
@@
More recently, the rise of Large Language Models (LLMs) for code generation has prompted their use in automating test generation [11, 10, 48, 22]. As unit test generators, LLMs tend to focus on components they infer as important—guided largely by the user’s prompt. The effectiveness of this process strongly depends on prompt quality [52], a challenge amplified in Rust due to its expressive and strict type system, which complicates the generation of valid code without precise context. To address this, researchers augment prompts with static analysis metadata and break down the generation task into smaller, focused sub-tasks. RUG [11] follows this approach, using ChatGPT-4 and static analysis to generate unit tests that capture complex trait relationships. While RUG’s high-quality test code improves coverage, its evaluation has not demonstrated discovery of memory safety bugs.
@@
deepSURF employs fuzzing to test the generated harnesses, shifting the burden of discovering bug-triggering conditions from the developer to the fuzzer and eliminating manual effort. To reduce false positives, it generates harnesses that invoke safe APIs capable of reaching unsafe code and configures the fuzzer to exclude crashes unrelated to memory corruption—a common issue in Rust fuzzing [39].
We evaluate deepSURF on 27 crates with real-world memory safety vulnerabilities and detect 26 bugs via fuzzing. Notably, six of these were previously-unknown and have been disclosed to the respective library authors. Our key contributions are summarized below.
@@
  • deepSURF Tool for Automatic Detection of Memory Safety Bugs. We implement deepSURF, a tool that combines LLM-augmented harness generation with fuzzing to detect memory safety bugs in Rust libraries. Compared to existing approaches, deepSURF shows significantly improved bug-finding capability, identifying 26 memory bugs—20 previously-known and six newly-discovered.
@@ 2 Background
Unsafe Rust: Superpowers with High Risk. Systems programming often requires low-level operations, such as interfacing with native C/C++ libraries—tasks that cannot be performed using “safe” Rust alone. To support these use cases, Rust provides unsafe Rust, which relaxes safety rules to give developers low-level control at the cost of reintroducing the risk of memory bugs.
@@
The effectiveness of LLM-based harness generation heavily depends on prompt quality [42, 53]. Yet, even with sufficient documentation and examples, LLMs often struggle to generate valid harnesses for complex targets like the Linux kernel [24] or Rust libraries [11, 48]. In Rust specifically, the combination of intricate trait relationships and a highly expressive type system poses major challenges. These complexities often prevent the automatic creation of valid harnesses, as accurately reasoning about such features is better suited to traditional static analysis techniques.
@@
Memory Safety Bugs in Rust Libraries. The use of unsafe Rust can introduce memory safety bugs due to incorrect handling of unsafe code [30, 49]. These bugs are similar to those found in other systems programming languages, including use-after-free, double-free, and buffer overflows. However, in Rust, not all memory violations constitute genuine memory bugs.
@@ 3 Challenges of Fuzzing Rust Libraries
Furthermore, although Rust confines potential memory vulnerabilities to unsafe code—unlike C, where all code is vulnerable—this introduces unique challenges in identifying and reaching that code. Unsafe code in Rust is governed by safety contracts that developers are expected to uphold. A memory violation caused by misusing a UAPI or incorrectly implementing an unsafe trait is typically attributed to the user for violating these contracts and does not constitute a true memory bug. Additionally, Rust fuzzers [36, 5] may classify panics and assertion failures as bugs, even when intentionally inserted by library developers. While such failures may indicate logical errors, they do not represent exploitable memory corruption and should be filtered out.
Generating harnesses to expose memory vulnerabilities in Rust requires overcoming several key challenges. Based on our analysis of real-world bugs in Rust libraries, in this section, we explain these key challenges and illustrate them with concrete examples.
1impl UnsafeSer for StructA {
2 fn process(&self) -> Vec<u8> {
3 unsafe { /* Unsafe operations */ }
4 }
5}
6
7impl SafeSer for StructA {
8 fn process(&self) -> Vec<u8> { ... }
9}
10
11pub unsafe fn unsafe_read(addr: usize) -> u8 {
12 /* Unsafe operations */
13}
14
15pub fn unsafe_ser<T: UnsafeSer>(data: T) -> Vec<u8> {
16 data.process()
17}
18
19pub fn safe_ser<T: SafeSer>(data: T) -> Vec<u8> {
20 data.process()
21}
Listing 1: Callee target resolution influenced by trait bounds.
@@
Even after identifying the exported APIs, determining which ones are URAPIs is non-trivial. Treating every API that reaches unsafe code as a URAPI can lead to false positives during fuzzing, especially if UAPIs are invoked directly in the harness. Avoiding this requires a more targeted control-flow analysis of the library. This analysis presents additional challenges. As shown in Listing 1, both unsafe_ser and safe_ser call process, but the method invoked depends on the trait bound—UnsafeSer or SafeSer. Resolving the correct callee requires analyzing the caller’s type context, which is further complicated by the use of dynamic dispatch [8]. Finally, beyond false positives from misuse of UAPIs, fuzzers may also classify crashes from developer panics and assertions as bugs [39]. To avoid this, the fuzzer must be configured to ignore non-memory corruption bugs.
@@
Challenge 3 (C3): Handling Generics Types. Rust supports generic data types, enabling reusability through type parameters. Generic types often have trait bounds specifying the behavior a type must implement. When an API expects a generic type, the harness must substitute it with a concrete type that satisfies the required trait bounds. This substitution is complex, as trait definitions may be scattered across the target or external libraries and some bounds may depend on supertraits [8]. Such complexity requires robust static analysis to extract complete trait information.
@@
3 fn read(&mut self, _: &mut [u8]) -> Result<usize> {
4 Ok(131313) // Return always the same big number
5 }
6}
7fn main() {
@@ 4 Ok(131313) // Return always the same big number
Challenge 4 (C4): Simulating Interaction with User-Defined Code. Rust libraries often allow users to define custom behavior through closures and traits. Closures are anonymous functions that can capture variables from their surrounding scope and are typically passed as arguments to enable dynamic behavior. Traits, by contrast, define required functions that describe a behavior types must implement. Users can create custom types with specific trait implementations, allowing libraries to adapt their functionality.
Our analysis of real-world memory safety bugs shows that library developers often fail to fully account for the range of behaviors that user-defined trait implementations or closures may exhibit—especially when interacting with unsafe code. For example, if a user-defined closure panics inside an unsafe block of the library, the Rust runtime will unwind the stack and invokes destructors for all live variables. If the unsafe code has duplicated ownership of any of these variables without proper safeguards, this can lead to double-free bug.
@@ 7 Ok(Window {
7 slab.remove(0); // Crashes due to HBOF
8}
@@
None of the existing tools configure the fuzzer to ignore non-memory corruption bugs, leading to false positives when crashes caused by panics or assertions are reported. Only RPG and RUG employ strategies to target unsafe code, but they may also directly fuzz UAPIs without preserving their safety invariants, which can also result in false positives. While all four tools support complex type arguments, some struggle when these types are nested within container types such as vectors. Although several approaches support generic types, none supports user-defined code simulation through custom trait implementations.
@@
The design of our tool is illustrated in Figure 1. deepSURF consists of three key components: Static Analysis, Harness Generation, and Dynamic Analysis. The tool processes a Rust library as input and attempts to detect memory safety vulnerabilities affecting it.
@@ 4.1 Unsafe Reachability Analysis
First of all, deepSURF needs to identify all URAPIs in the provided Rust library, since these are the functions that can potentially trigger the execution of unsafe code. This process consists of two steps: (a) detecting all locations of unsafe code within a library and (b) finding publicly accessible safe functions that can reach those locations.

### v2

# Accepted to IEEE Symposium on Security and Privacy 2026 deepSURF: Detecting Memory Safety Vulnerabilities in Rust Through Fuzzing LLM-Augmented Harnesses

@@ Abstract
To address these limitations, we present deepSURF, a tool that integrates static analysis with Large Language Model (LLM)-guided fuzzing harness generation to effectively identify memory safety vulnerabilities in Rust libraries, specifically targeting unsafe code. deepSURF introduces a novel approach for handling generics by substituting them with custom types and generating tailored implementations for the required traits, enabling the fuzzer to simulate user-defined behaviors within the fuzzed library. Additionally, deepSURF employs LLMs to augment fuzzing harnesses dynamically, facilitating exploration of complex API interactions and significantly increasing the likelihood of exposing memory safety vulnerabilities. We evaluated deepSURF on 63 real-world Rust crates, successfully rediscovering 30 known memory safety bugs and uncovering 12 previously-unknown vulnerabilities (out of which 11 have been assigned RustSec IDs and 3 have been patched), demonstrating clear improvements over state-of-the-art tools.
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ 1 Introduction
Rust’s focus on prioritizing memory safety while maintaining high performance makes it a compelling competitor in the systems programming space. Unlike traditional low-level languages like C/C++, which rely on manual memory management and are prone to common memory safety bugs such as buffer overflows, Rust’s unique set of safety rules eliminates these risks as early as during compilation. This feature has enabled its adoption in mainstream operating systems [23, 43], while its public registry of over 199k crates [^1] In Rust, a crate is a unit of code that can be a library or a binary. highlights its growing popularity  [14].
@@
Unsafe Rust undermines the language’s memory safety guarantees, introducing memory safety vulnerabilities. Microsoft and Google report that over 70% of critical bugs stem from memory safety issues [42, 56], leading governmental agencies such as CISA and NSA to advocate for memory-safe programming [16, 31]. To date, around 27% of Rust bugs in the RustSec Advisory Database [55] are related to memory corruption, resulting from incorrect use of unsafe Rust [35, 63]. This significant percentage has motivated research into static[8, 36, 41] and dynamic [15, 5, 6] analysis tools to detect such issues.
@@
Static analyzers such as Rudra [8] and Yuga [36] aim to identify certain bug patterns in Rust source code. Although these tools successfully detect memory safety bugs, significant human involvement is required to filter their output (with high false positive rates) and to create proof-of-concept (PoC) test cases to confirm each bug. In contrast, dynamic analysis approaches, such as fuzzing, due to their context-sensitive analysis tend to have lower false positives and can uncover a broader range of bugs [6, 5, 47]. Although fuzzers are effective at detecting memory safety violations, their utility is often limited in the Rust ecosystem, where most code is provided as libraries rather than standalone binaries that can be directly tested by the fuzzer [26]. Consequently, before fuzzing Rust libraries, it is necessary to generate appropriate harnesses that encapsulate the library’s functionality and convert the fuzzer’s input into well-typed, valid data types expected by the target functions. Automated harness and test generation are well-studied in other languages: Randoop [39] uses execution feedback to generate Java unit tests, and Pynguin [32] extends this approach to Python.
@@
Harness generation in Rust is tedious and time-consuming due to its complex type system and syntax. To ease this burden, several tools use static analysis to automate harness generation. However, despite improving test coverage, these tools struggle to detect memory safety bugs. RULF [26] and FRIES [66] lack support for traits, and RuMono [69] does not handle closures—both crucial for modeling buggy user-defined behavior. RPG [64] targets all unsafe code, including explicitly unsafe APIs, resulting in false positives. Additionally, these tools use heuristics and API dependency analysis to generate API call sequences. While these sequences can be syntactically valid, their API calls are often not semantically related and fail to reflect realistic usage patterns. Due to these limitations, although these tools have discovered bugs in various Rust crates, none of the reported issues have involved memory safety.
@@
More recently, the rise of Large Language Models (LLMs) for code generation has prompted their use in automating test generation [12, 11, 60, 25, 59] and fuzzing [61]. As unit test generators, LLMs tend to focus on components they infer as important—guided largely by the user’s prompt. The effectiveness of this process strongly depends on prompt quality [67], a challenge amplified in Rust due to its expressive and strict type system, which complicates the generation of valid code without precise context. To address this, researchers augment prompts with static analysis metadata and break down the generation task into smaller, focused sub-tasks. RUG [12] follows this approach, using GPT and static analysis to generate unit tests that capture complex trait relationships. While RUG’s high-quality test code improves coverage, its evaluation has not demonstrated discovery of memory safety bugs.
@@
deepSURFemploys fuzzing to test the generated harnesses, shifting the burden of discovering bug-triggering conditions from the developer to the fuzzer and eliminating manual effort. To reduce false positives, it generates harnesses that invoke safe APIs capable of reaching unsafe code and configures the fuzzer to exclude crashes unrelated to memory corruption [47].
We evaluate deepSURF on 63 real-world Rust crates and detect 42 memory safety bugs via fuzzing. Notably, 12 of these were previously-unknown and have been disclosed to the respective library authors. Our key contributions are summarized below.
@@
  • deepSURF Tool for Automatic Detection of Memory Safety Bugs. We implement deepSURF, a tool that combines LLM-augmented harness generation with fuzzing to detect memory safety bugs in Rust libraries. Compared to existing approaches, deepSURF shows significantly improved bug-finding capability, identifying 42 memory bugs—30 previously-known and 12 newly-discovered. We release the source code of deepSURF publicly [40].
@@ 2 Background
Rust Language Features. In addition to its ownership and lifetime rules, Rust introduces several language features that are central to its programming model.
Traits. Traits in Rust are similar to interfaces in Java: they specify a set of required methods (hereafter, trait functions) that a type must implement to provide a given behavior. For example, a type implementing the Iterator trait must define how to produce the next element in a sequence by implementing the next trait function. Traits may also define associated types—trait-local type parameters fixed per implementation—clarifying trait’s input/output relationships. Traits enable reusability and flexibility by letting code operate on any type that implements the trait.
Closures. Closures in Rust are anonymous functions that can capture variables from their surrounding scope. They are often passed as arguments to functions, enabling flexible and dynamic behavior (e.g., passing a closure to specify a custom condition for filtering elements of a data structure).
Panics. A panic in Rust represents an unrecoverable program state (e.g., out-of-bounds array access, unwrapping the None type or requesting more memory than is available). When a panic occurs, the Rust runtime, by default, begins unwinding the stack and terminates the program’s execution in a controlled way. Panics can be triggered explicitly with the panic! macro to indicate violated preconditions or failed assertions. For example, APIs that allow indexing into queue-like structures often begin by asserting that the index is within bounds; otherwise, they panic. A triggered panic may not always correspond to a bug in the code, but it often serves as a deliberate guard preventing more severe failures.
Unsafe Rust: Superpowers with High Risk. Systems programming often requires low-level operations, such as interfacing with native C/C++ libraries—tasks that cannot be performed using “safe” Rust alone. To support these use cases, Rust provides unsafe Rust, which relaxes safety rules to give developers low-level control at the cost of reintroducing the risk of memory safety bugs.
@@
The effectiveness of LLM-based harness generation heavily depends on prompt quality [50, 68]. Yet, even with sufficient documentation and examples, LLMs often struggle to generate valid harnesses for complex targets like the Linux kernel [27, 65] or Rust libraries [12, 60]. In Rust specifically, the combination of intricate trait relationships and a highly expressive type system poses major challenges. These complexities often prevent the automatic creation of valid harnesses, as accurately reasoning about such features is better suited to traditional static analysis techniques.
@@
Listing 1 shows concrete examples of Rust functions, with each function labeled by its applicable categories.
Validity of Rust Library Fuzzing Findings. The fuzzing of a Rust library harness may terminate prematurely due to runtime failures such as Rust panics, or memory errors (e.g., segmentation faults). In the context of fuzzing, we refer to these events as crashes. Not all crashes indicate a library bug, so we classify them as follows.
  • True positives: Crashes caused by a developer-introduced defect in the library that can be triggered under correct (contract-respecting) use of the library. In other words, the harness follows the library according to its documentation (in terms of types, preconditions, invariants), yet erroneous or unsafe behavior occurs.
  • False positives: Crashes that are not caused by library bugs, including (a) intended behavior when API contracts are violated (e.g., calling an API with arguments documented to cause a panic or memory error), and (b) panics or memory errors in the harness code itself (e.g., invalid type conversions, intentional panics, or introduction of incorrect unsafe code).
In Appendix B, we provide additional justifications and empirical data regarding this distinction.
1impl UnsafeSer for StructA {
2 fn process(&self) -> Vec<u8>{ unsafe { ... } } // SF, UEF
3}
4impl SafeSer for StructA {
5 fn process(&self) -> Vec<u8>{ ... } // SF
6}
7fn do_safe_ser<T: SafeSer>(dat: T) -> Vec<u8>{ // SF
8 dat.process()
9}
10fn do_unsafe_ser<T: UnsafeSer>(dat: T) -> Vec<u8>{ // SF, URF
11 dat.process()
12}
13pub unsafe fn unsafe_read(addr: usize) -> u8{ // UF, UAPI
14 /* Unsafe operations */
15}
16pub fn safe_ser(sA: StructA) -> Vec<u8>{ // SF
17 do_safe_ser(sA)
18}
19pub fn unsafe_ser(sA: StructA) -> Vec<u8>{ // SF, URF, URAPI
20 do_unsafe_ser(sA)
21}
Listing 1: Callee target resolution influenced by trait bounds.
Memory Safety Bugs in Rust Libraries. The use of unsafe Rust can introduce memory safety bugs due to incorrect handling of unsafe code [35, 63]. These bugs are similar to those found in other systems programming languages, including use-after-free, double-free, and buffer overflow violations. However, in Rust, not all memory violations constitute genuine memory bugs.
@@ 3 Challenges of Fuzzing Rust Libraries
Furthermore, although Rust confines potential memory vulnerabilities to unsafe code—unlike C, where all code is vulnerable—this introduces unique challenges in identifying and reaching that code. Unsafe code in Rust is governed by safety contracts that developers are expected to uphold. A memory violation caused by misusing a UAPI or incorrectly implementing an unsafe trait is typically attributed to the user for violating these contracts and does not constitute a true memory bug. Moreover, Rust existing fuzzers [44, 5] classify assertion failures and panics as crashes, even when they are intentionally inserted by library developers to prevent incorrect API use [47]. Such crashes are often false positives and increase the amount of manual triage required to analyze the fuzzer’s findings; therefore, the fuzzer should be configured to automatically filter these crashes out.
Generating harnesses to expose memory corruption vulnerabilities in Rust requires overcoming several key challenges. Based on our analysis of real-world bugs in Rust libraries, in this section, we explain these key challenges and illustrate them with concrete examples.
@@
Even after identifying the exported APIs, determining which ones are URAPIs is non-trivial. Treating every API that reaches unsafe code as a URAPI can lead to false positives during fuzzing, especially if UAPIs are invoked directly in the harness. Avoiding this requires a more targeted control-flow analysis of the library. This analysis presents additional challenges. As shown in Listing 1, both do_unsafe_ser and do_safe_ser call process, but the method invoked depends on the trait bound—UnsafeSer or SafeSer. Resolving the correct callee requires analyzing the caller’s type context, which can be further complicated when using dynamic dispatch [9].
Finally, beyond false positives from misuse of UAPIs, fuzzers may also classify assertion failures and panics as crashes, even when inserted intentionally by library developers to enforce invariants and API contracts [47]. While some of these failures may be related to genuine bugs, their majority are either false positives or bugs that are not exploitable memory corruptions. Because our focus is on detecting memory corruption bugs while minimizing the manual effort required to analyze fuzzer crashes, we should filter these failures out. We acknowledge that this design choice may lead to false negatives, and we evaluate this aspect in Appendix B.
@@
Challenge 3 (C3): Handling Generic Types. Rust supports generic data types, enabling reusability through type parameters. Generic types often have trait bounds specifying the behavior a type must implement. When an API expects a generic type, the harness must substitute it with a concrete type that satisfies the required trait bounds. This substitution is complex, as trait definitions may be scattered across the target or external libraries and some bounds may depend on supertraits [9]. Such complexity requires robust static analysis to extract complete trait information.
@@
3 // Return always the same incorrect number
4 fn read(&mut self, _: &mut [u8]) -> Result<usize> {Ok(999)}
5}
6fn main() {
@@
Challenge 4 (C4): Simulating Interaction with User-Defined Code. Rust libraries often allow users to define custom behavior through closures and traits. User-defined closures can be passed as API arguments to enable custom dynamic behavior. In addition, users can create custom types with specific trait implementations, allowing libraries to adapt their functionality.
Our analysis of real-world memory safety bugs shows that library developers often fail to fully account for the range of behaviors that user-defined trait implementations or closures may exhibit—especially when interacting with unsafe code. For example, if a user-defined closure panics inside an unsafe block of the library, the Rust runtime will unwind the stack and invoke destructors for all live variables. If the unsafe code has duplicated ownership of any of these variables without proper safeguards, this can lead to a double-free bug.
@@ 7 Ok(Window {
6 slab.remove(0); // Crashes due to heap buffer overflow
7}
@@
None of the existing tools configure the fuzzer to ignore crashes unrelated to genuine bugs, leading to false positives when panics or assertion failures are triggered due to incorrect API invocation. Only RPG and RUG employ strategies to target unsafe code, but they may also directly fuzz UAPIs without preserving their safety contracts, which can also result in false positives. While all five tools support complex type arguments, some struggle when these types are nested within container types such as vectors. Although several approaches support generic types, none supports user-defined code simulation through custom trait implementations.
@@
The design of our tool is illustrated in Figure 1. deepSURF consists of three key components: Static Analysis, Harness Generation, and Dynamic Analysis. The tool processes a Rust library as input and attempts to automatically detect memory safety vulnerabilities affecting it.
@@ 4.1 Unsafe Reachability Analysis
First of all, deepSURF needs to identify all URAPIs in the provided Rust library, since these are the functions that can potentially trigger the execution of unsafe code. This process consists of two steps: (a) detecting all locations of unsafe code within a library and (b) finding publicly accessible safe functions that can reach these locations.